OpenStudy (anonymous):

The sum of two numbers, x and y, is 12. The difference of x and two times y is 6. What are the values of x and y? A.) x=8, y=4 B.) x=10, y=2 C.) x=18, y=-6 D.) x=20, y=-8 Please Help

OpenStudy (aripotta):

you have to set up a system of equations. "the sum of two numbers is 12" means x + y = 12. "the difference of x and two times y is 6" means x - 2y = 6. so those are our two equations x + y = 12 x - 2y = 6 there are different ways to solving this kind of problem. i think i'd go with the elimination method...

OpenStudy (aripotta):

to do that, i'm going to multiply the whole second equation by -1, changing all the signs in the equation so -1(x - 2y = 6) is -x + 2y = -6

OpenStudy (aripotta):

then i'm going to combine the equations and eventually solve for y x + y = 12 + -x + 2y = -6 3y = 6 y = 2

OpenStudy (anonymous):

so the answer is ?

OpenStudy (aripotta):

and the only choice with y as 2, is b

OpenStudy (anonymous):

thank you so much i have more questions also

OpenStudy (aripotta):

i will help you only if you try it out yourself first :)

OpenStudy (anonymous):

Consider this system of linear equations: y − x = -13 -4x + 3y = -51 The solution of the system of equations is A.) 112/5, 47/5 B.)-12, 1 C.) -12, -1 D.) 12,-1

OpenStudy (anonymous):

Y - x = -13

OpenStudy (aripotta):

do you have any ideas?

OpenStudy (anonymous):

none at all like im very horrible at math i do not understand any of this what so ever

OpenStudy (aripotta):

ok, so for the sake of teaching, i'm going to use a different method this time; the substitution method...

OpenStudy (aripotta):

y - x = -13 -4x + 3y = -51 first, i'm going to use the first equation to solve for y. once i add x to both sides, y = -13 + x. now that i know what y equals in terms of x, i'm going to substitute that into the second equation. -4x + 3y = -51 -4x + 3(-13 + x) = -51 distribute the 3 -4x - 39 + 3x = -51 combine like terms -x - 39 = -51 add 39 to both sides -x = -12 divide both sides by -1 to make -x positive x = 12 and there's no reason to do any further math, because the only choice with x as 12 is d

OpenStudy (anonymous):

so the answer is D?

OpenStudy (aripotta):

yes

OpenStudy (aripotta):

and if you have any more questions, i'm not going to do everything for you next time

OpenStudy (aripotta):

are you understanding what i'm doing?

OpenStudy (anonymous):

no not at all like math is so hard i dont understand math at all its the hardest subject ive ever taken ive already failed algebra 1 once now im just doing credit recovery so i can get into my right grade which is 11th

OpenStudy (aripotta):

do you have any questions about what i did to get the answer?

OpenStudy (aripotta):

asking questions about stuff you don't understand is a key element to learning

OpenStudy (aripotta):

if you don't ask questions, how could i possibly know what you don't understand?

OpenStudy (anonymous):

umm not really cause if you explained it i wouldnt get it

OpenStudy (anonymous):

its just the WHOLE algebra 1 subject i dont understand anything

OpenStudy (aripotta):

regarding this problem, what don't you get? i explained every step. does my explanation help at all?

OpenStudy (anonymous):

not really because i dont understand it

OpenStudy (aripotta):

my goal right now is to help you understand. look back to my explanation and tell me any terms you don't know

OpenStudy (anonymous):

like i know what the terms mean but i just dont know like where to start and how to finish it

OpenStudy (anonymous):

12x= 54-6y? -17x= -62 -6y ?

OpenStudy (aripotta):

when you have a system of equations, which is what these problems are, you can solve them by the elimination method (which i did for the first problem), the substitution method (i did for the second), and by using matrices (which aren't really relevant right now, unless you're in honors algebra)

OpenStudy (anonymous):

well i dont think ill be in honors anytime soon

OpenStudy (aripotta):

is that the next problem?

OpenStudy (anonymous):

Solve this system of linear equations. Separate the x- and y-values with a comma. 12x= 54-6y? -17x= -62 -6y ?

OpenStudy (aripotta):

ok

OpenStudy (aripotta):

look at the y's. what do you notice about them?

OpenStudy (anonymous):

they both have a -6 infront of them

OpenStudy (aripotta):

exactly. to use the elimination method, we want to multiply the entire second equation by -1, so that all the signs are changed. that will give us positive 6y. and when we combine the equations, the y will be cancelled out. so -1(-17x= -62 -6y) is 17x = 62 + 6y

OpenStudy (anonymous):

do i need a calculator?

OpenStudy (aripotta):

now we have our equations as 12x = 54 - 6y 17x = 62 + 6y when we combine them, we'll get 29x = 116

OpenStudy (aripotta):

a calculator would be helpful :)

OpenStudy (aripotta):

see how we completely got rid of the y? now we can solve for x

OpenStudy (anonymous):

alright i got a calculator

OpenStudy (anonymous):

well my class ends at 2:05 pm and its 1:45 pm now

OpenStudy (aripotta):

we're almost done

OpenStudy (aripotta):

so 29x = 116. divide both sides by 29 and you get x = 4

OpenStudy (anonymous):

alright well im on 4 out of 15 questions

OpenStudy (aripotta):

oh.

OpenStudy (aripotta):

now that we know what x equals, you can substitute it into either equation to solve for y

OpenStudy (aripotta):

12x = 54 - 6y 12(4) = 54 - 6y 48 = 54 - 6y -6 = -6y y = 1

OpenStudy (aripotta):

so your answer is 4, 1

OpenStudy (anonymous):

-17x = -95 -10y 12x =50 -10y

OpenStudy (aripotta):

maybe i can help you more after school?

OpenStudy (anonymous):

yeah probably cause i have internet at home

OpenStudy (aripotta):

see, i don't wanna just work at it myself. i want you to understand what to do

OpenStudy (aripotta):

i don't think i'm helping you at all, am i?

OpenStudy (anonymous):

ah i understand well thank you for helping me with the answers and i get out of school at 3:40 PM and i normally get home like around 4:20 something

OpenStudy (aripotta):

alright. just tag me in any questions :)

OpenStudy (anonymous):

how do i tag ?

OpenStudy (aripotta):

@ and then my name so @aripotta

OpenStudy (anonymous):

alright thanks :)

OpenStudy (aripotta):

\(\Large \boldsymbol{ \rlap{\color{#008AFF}{\hspace{2.5pt}n}\color{#0073FF}{\hspace{0.34pt}o} \hspace{3pt}\color{#005CFF}{p} \color{#0045FF}{\hspace{0.9pt}r} \color{#0017FF}{o} \color{#0200FF}{\hspace{1.7pt}b} \color{#3000FF}{\hspace{1.5pt}l} \color{#4700FF}{\hspace{0.4pt}e} \color{#7500FF}{\hspace{0.9pt}m} \color{#8C00FF}{\hspace{0.5pt}o}~\LARGE\color{#D100FF}{\ddot\smile}}{\hspace{1.5pt} no~p\hspace{0.3pt}r\hspace{0.4pt}o\hspace{1pt}b\hspace{1.5pt}l\hspace{0.4pt}e\hspace{0.4pt}m\hspace{0.4pt}o}}\)